Summary of the GMVA Session, Sep. 29 - Oct. 4 2016 ================================================== Participating stations: Effelsberg, Pico, PdBure, Yebes, Onsala, Metsahovi, VLBA, GBT, 3xKVN, (Haystack) General Summary: ================ t162a is a wide band (single IF, 500 MHz) recording test within Europe (no IRAM). be074 is a fringe test between VLBA (FD,KP,MK only) and ATCA, ATCA supports only 1 Gbps, 2 IFs, linear pol. The intitial fringetest (f162a) showed fringes for Effelsberg, PdBure, Yebes, Onsala, but not for Pico and Metsahovi. Weather was very poor in Effelsberg, Onsala and Metsahovi, and better in Pico, PdBure and Yebes. The VLBA had also bad weather at some stations, VLBA_KP dropped off because of a VME problem. Haystack participated as test in some scans on 3C84 (c162b). But one polarisation was broken and the weather was bad. Pico and PdB did not observe 3C84 (MH001A in c162b), because of rejection by TAC. Detailed logs: t161a - Wide band fringe test Participating: Effelsberg (DBBC), Onsala, Yebes stations should use V16 of the DBBC firmware Pico has problems with V16, will stay on V14, don't participate Uwe says we should same setup as in spring (t161cef.prc) Yebes: cannot interleave polarisations, just go for one polarisation followed by the other, so this sequence: vsi1=a05,a06,a07,a08,a09,a10,a11,a12,b05,b06,b07,b08,b09,b10,b11,b12 Onsala: recording test yesterday with PFB V16, sampler statistics and the auto-correlation bandpass looked reasonable good, will use V16 Onsala confirms vsi1 list from above. This is consistent with the vsi1 list created by FS 9.11.16. be074 - ATCA -VLBA, 3mm test VLBA: recording should be on the same modules used for this GMVA session ATCA should e-transfer data to Bonn if possible VLBA log: 2027-2231 %MK Focus drive faults. However, focus values are close to commanded position and aren't changing all that much between scans, so data should still be fine. 2215 *KP raining. ATCA: IF1 recording to DAS1 to cavsi1, IF2 recording to DAS2 to cavsi2. The statistics on channel 1 DAS2 look strange, but the DAS seems to be working as required. Pcal on 1055+018 at 21:31:13. The phase stability after this time seemed good enough to include CA03 in the tied array as well, giving us a four element tied array. We are not including CA02 because it has unequal power on its polarisations and may cause problems in the array addition. The tied array went from 145 to 1345 at 21:38:02. f162a - European fringe test Effelsberg: start in good weather, clear, use RDBE, Tsys~190K, connection to 2nd Horn (Weather Horn) is unstable, may not be able to use it for pointing. Pico: We have join the f162a schedule on scan number 4 under bad weather conditions, thick clouds. Tsys is about 800k !! at 3mm and 1749+096 is not visible for now. Maser-GPS is about 1.3E-13. PdB: NOEMA started with some delay. We have missed the first fringe test scan but have started to transfer the second to Grenoble (scan 9). We are observing under clear sky with 6 phased antennas. Our reference antenna pad position is N11. We observe with Integer-Frequencies (10 kHz OFF, no comb). Onsala: Our DBBC IF B had proper IF level. While, the total power output from power commands were quite high since 17:00 UT. No idea why. Metsa: Extremely bad weather last evening and night, real fall storm and rain. Forecast promises that this should be over soon. Yebes: No apparent problems. Good weather Correlator: Fringes between Eb_RDBE-PB and Eb_RDBE-Ys, Ys has only 1 pol (LCP), and Eb_RDBE-Onsa (peculiar fringe rate, weather?) Cannot find fringes to PV, reason unclear. c162a Pico: Now we have started partA with better weather conditions, Tsys about 130k . An issue with temperature stabilization of the dish requires frequent focus checks (once every 1.5hrs) Effelsberg: Sep 30, 12 UT overcast, pointing ok 18 UT ok weather poor, often use 13 GHz (P22mm) RX for pointing Yebes: No apparent problems. Good weather Onsala: Problems with total power overflow/underflow in PFB subbands at Onsala. Sampler statistics poor, restart DBBC (without firmware reload), problem temporarily goes away. KVN_TN: failing hexapod in sub-reflector, did not observe VLBA: 0354-0449 *PT,MK removed for USNO observation 0535-2023 *KP VME is not communicating 0755-0848 *PT,MK out for USNO re-observation 1435 KP site tech reports VME power supply is dead KP antenna has been stowed and data is unuseable. c162b Pico: We have resumed the VLBI recording on partB at 10:15 UT We tested also the phase stability of the 1st IF from the receiver cabin to the DBBC, the pity is that we couldn't inject in front of the receiver due to the broken harmonic generator, but that should be no problem since the receiver was properly locked. Effelsbg: raining, Oct. 1, 0-10 UT Oct 1, 15h30 clearing up, do pointing/focus on Venus, resume VLBI 16h00 PdB: bad weather, no data Yebes: No apparent problems. Good weather Hays: we lost power to the second IF fiber, leaving us with only a single polarization channel, The weather was also fairly poor (intermittent light rain), although we were able to see an SiO maser source for pointing. Given these factors, we decided to record only a single scan on each of the program targets for fringe test purposes. VLBA: 0257-0323 *PT Out for USNO observation 0327-1652 *KP out. VME power supply dead. KVN_YS: a fraction of data lost due to failure on disk storage server (only 58 scans e-transferred to Bonn) c162c PdB: weather quite bad, resume 23h local time (Oct 1st?) with new reference antenna on Pad N17 Problem subreflector A6 - need to change the reference antenna Reference antenna 7 on N17 A software problem may cause the lack of fringes. Oct. 1 17-20 UT bad weather Oct. 1 20h0 -> 02h30 UT ok Oct. 2 03h00 -> 04h30 UT bad weather Oct. 2 05h30 -> 10h00 UT OK Oct. 2,10h30 : change back from Reference antennta A7 to A6 on N11. fringes may come back Oct. 2 10h30 -> end done Pico: Oct 2, weather very good Yebes: Detected power at the DBBC dropped to 0 between 00:15 and 6:15 on Oct 2. A new PFB control software version (300916) was loaded at 06:15 and the power levels were correct after that until the end of the experiment. Good weather. Effelsbg: Oct.2, 7 UT, last night bad, rain, focus was off for 2nd half of night Oct.2, 14 UT cloudy, pointing difficult Oct.3, 8-12 UT, rain, after 12 UT rain stopped, ptg. ok Onsa: ok, underflow/overflow in PFB subbands has not been seen for two days GBT: GBT starts Oct.1st, weather was bad, no rain, but heavy clouds with zenith opacity estimated at > 0.5 for the first several hours. Tsys 150-250 K 03:30 UT - 3C454.3, conditions were better, but still not very good, zenith opacity 0.25 or so, and Tsys 120-150K. next evening - The weather is slightly better. Opacity about 0.24; Tsys still 120-150K. peak scans are done in linear mode, cal and VLBI scans in circular mode (1/4 plate) VLBA: 1657-1833 *KP not observing due to dead VME power supply. 2006-2301 *OV powerloss, no pointing Oct2: 0255-0349 *PT,MK out for daily USNO observation 0623-1634 *KP communications down. VME may have died again Oct3: 0225-0319 *PT,MK out for daily USNO observation. c162d Pico: We have started c162d @94GHz in good weather conditions,Tsys 130k. NRAO530 couldn't be recorded, we were late to the schedule. Ef: we started c162d , ptg and focus on 3c454.3, then on NRAO530. We can see NRAO530. Weather clear just before start, but at 18h30 to end light rain and thick clouds. Yebes: Yebes has started on time but we can discard data until 21:30 local time (19:30 UT). We have been fixing issues with the receiver. Tsys now is 370 K approx. VLBA: Ants Scheduled: NL FD KP OV MK 2110 %MK overcast, site tech reports rain. Detailed VLBA Logs: =================== VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C162A Observer: Krichbaum, T. Project type: OTHER Obs filename: c162acrd.* Date/Day: 2016SEP29/273 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is A. Sowinski. UT1-UTC -0.28; gps delta and vclock okay. 2327 Begin at NL,FD,LA. 2342 PT joins. 2349-0719 *PT raining. 2357 KP joins. 0000 Date is 2016SEP30/274. 0012 OV,BR join. 0242 MK joins. 0242-0354 *MK rotation encoder coupling slipped, weak or no FRM 100 72 fringes (C138908). 0354-0449 *PT,MK removed for USNO observation N6274. SCH 200 110 0449 %PT,MK may experience phase and delay jumps when rejoining the observation. 0449-0755 *MK rotation encoder coupling slipped, weak or no FRM 100 186 fringes (C138908). 0500-0501 *MK Mark5C bad-record-rate. REC 100 1 0535-2023 *KP VME is not communicating. 0543 KP reset power on VME. 0545 New operator is Alan Kerr 0603 KP VME chatterbox reset again. 0630 KP VME chatterbox reset (third try). still did not come back up. assuming that the VME is not pointing the antenna any more. 0755-0848 *PT,MK out for USNO re-observation: N6274B SCH 200 106 0848 %PT,MK may experience phase and delay jumps when rejoining the observation. 0848-2023 *MK rotation encoder coupling slipped, weak or no FRM 100 695 fringes (C138908). 1344 New Operator: Paul Padilla 1435 KP site tech reports VME power supply is dead KP antenna has been stowed and data is unuseable. 1520-1522 *LA autostowed, power glitch? COM 100 2 2023 End. Downtime Summary: Total downtime : 1172 min Percentage downtime of observing: 11.7% Average downtime per hour : 7.0 min Total scheduled observing time (# Antennas): 10048 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard. VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C162B Observer: Krichbaum, T. Project type: OTHER Obs filename: c162bcrd.* Date/Day: 2016OCT01/275 Ants Scheduled: NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is A. Sowinski. UT1-UTC -0.28; gps delta and vclock okay. 0227 Begin at NL,BR. 0257 LA,PT join. 0257-0323 *PT Out for USNO observation N6275. SCH 100 26 0312 FD joins. 0327 OV,KP join. 0327-1652 *KP out. VME power supply dead. HWC 100 805 0337-0346 *LA raining. 0545 New operator is Alan Kerr 0641 Begin at MK. 0844-1135 %MK precipitation 1156-1205 *OV power outage. COM 100 9 1209 %NL raining. 1247-1258 *OV power outage. COM 100 11 1356 New operator is Betty Ragan. 1517 %MK raining. 1650 End at KP, OV, MK. 1652 End. Downtime Summary: Total downtime : 851 min Percentage downtime of observing: 12.3% Average downtime per hour : 7.4 min Total scheduled observing time (# Antennas): 6920 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard. VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C162C Observer: Krichbaum, T. Project type: OTHER Obs filename: c162ccrd.* Date/Day: 2016OCT01/275 Ants Scheduled: SC HN NL FD LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Betty Ragan. Vclock, gps delta, UT1-UTC all check OK. 1657 Begin at all sites except GB,MK 1657-1833 *KP not observing due to dead VME power supply. HWC 100 96 1820 Begin at GB. 1820 %GB is showing zero tsys. There is no obvious reason for that on our end of things. 1833 %KP power supply has been replaced, and it's back in. We are seeing an issue where the station clock and the GPS disagree, with the amount of disagreement fluctuating rapidly back and forth from about -2 to +4 seconds. The station clock actually appears to be keeping steady time, however, the gps delta is good, and vclock is showing a reasonable Clk-UTC time difference, so this is probably not an actual timing problem. (C138882) 2012 Begin at MK. 2006-2301 *OV lost power, is on generator. Unable to COM 100 175 point the antenna. 2016-1044 %MK raining. 2222-2223 *KP autostowed. Power glitch. COM 100 1 2329-2331 *LA autostowed. Power glitch. COM 100 2 0000 Date is 2016OCT02/276. 0145 New operator is Alan Kerr 0255-0349 *PT,MK out for daily USNO observation. SCH 200 108 0304-0316 *GB still a Tsys of zero on all channels. Device TIME 100 12 Browser now showing no communications to the m401 module. rebooting everything: restarting executor, rebooting both RDBEs, reloading obs. no change. m401 still not communicating. 0325 GB operator believes that the "zero Tsys" is due to a difference in hardware between the VLBA and GB systems. 0623-1634 *KP communications down. VME may have died again. HWC 100 611 0636-0640 *MK autostow. power glitch. repointed. COM 100 4 1358 New operator is Betty Ragan. 1633 KP is back up. Site tech replaced the hard drive. 1639-1916 *KP is still having issues. The VME appears to be having some communications problems. Site tech is taking it back to work on it some more. 1708 %NL raining. 1917-1924 *KP VME not sending back monitor data. Rebooted it. HWC 100 7 2132 %KP raining. 2315 %OV windy. 0000 Date is 2016OCT03/277. 0104-0636 %MK raining. 0145 New operator is Alan Kerr 0225-0319 *PT,MK out for daily USNO observation. SCH 200 108 0554-0619 %PT precipitation 1010-1109 %LA precipitation 1135 %OV windy 1351 New operator is Betty Ragan. 1503-1509 *PT has timing error error on DBE-1. TIME 100 6 1557-1614 *KP VME partially stopped communicating again and had HWC 100 17 to be rebooted. Didn't come up after the site techs reset it, or after using the chatterbox. Site techs to power it off and back on again. Antenna stowed while the VME was down and had to be repointed. 1722 End. Downtime Summary: Total downtime : 1147 min Percentage downtime of observing: 3.9% Average downtime per hour : 2.4 min Total scheduled observing time (# Antennas): 29050 min (10) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard. VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C162D Observer: Krichbaum, T. Project type: OTHER Obs filename: c162dcrd.* Date/Day: 2016OCT03/277 Ants Scheduled: NL FD KP OV MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Betty Ragan. Vclock, gps delta, UT1-UTC all check OK. 1927 Begin. 1927 %PT windy. 2110 %MK overcast, site tech reports rain. 2128 New Operator: Paul Padilla 0000 Date is 2016OCT04/278. 0015 End. Downtime Summary: Total downtime : 0 min Percentage downtime of observing: 0.0% Average downtime per hour : 0.0 min Total scheduled observing time (# Antennas): 1440 min (5) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.